EH001

PI: Hong
Title: 1156+295

Session: Feb 1997 (session 1)
Frequency: 6 cm
Mode: B
Obs time (UT): EU 19:30(21/02)-07:00(22/02)
Stations: EbHhMcNtOnShSmTrUrWb+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A. Cattani at Sat Feb 22 02:24:40 EST 1997) Success

Onsala (Antonis Polatidis at Sun Feb 23 09:04:28 EST 1997) Minor failures
Due to high winds the antenna was stowed until UT 0247. No known problems with the recorded data.

Effelsberg (dgraham at Mon Feb 24 03:29:45 EST 1997) Failure
Telescope drive failed from 0000 until end of experiment

Westerbork (Tony Foley at Mon Feb 24 10:54:14 EST 1997) Minor failures
One restart (maybe phase jump) at 06:30 UT

HartRAO (Jonathan Quick at Mon Feb 24 21:14:12 EST 1997) Minor failures
Read heads failing, write head status unknown

Shanghai (Huang Xinyong at Wed Mar 5 04:57:52 EST 1997) Minor failures
Clock offset = 0.10 msec. Tape vacuum failed between 19:35:10 - 19:37:04. Weather fine.

Jodrell (Alastair Gunn at Wed Mar 5 09:06:00 EST 1997) Minor failures
FS re-started (new log) due to problems with vacuum at start of experiment. Data should be good from 20:41 (number 5) scan. FS crashed and re-started at 02:16 UT. No other known problems. Tapes EVN00043,B284 and MPIR1221,F37B sent to Bonn via Woodcroft on 24/02/97.

Cambridge (Alastair Gunn at Wed Mar 5 09:10:15 EST 1997) Success
No known problems. Tapes NAO00004,B6F3 and NGS00390,32EF sent to Bonn via Woodcroft on 24/02/97.

Calibration (Fredrik T Rantakyro at Thu May 8 11:35:22 METDST 1997) Success
The ANTAB file for the experiment has been placed on VLBEER in the Feb97 directory with the name eh1.antab. This file contains the preliminary calibration results and may be updated when I receive updated information from the telescopes and the results from the datareduction of the calibration experiments.

MPIfR (Scott Aaron at Fri Mar 20 10:55:37 MET 1998) Minor failures
Correlation completed. Usual tracking problems with Torun. Frequent problems with the Jodrell tapes too. No other problems reported.